# sample.bib

### v1

@@ Abstract
With the development of artificial intelligence and breakthroughs in deep learning, large-scale Foundation Models (FMs), such as GPT, CLIP, etc., have achieved remarkable results in many fields including natural language processing and computer vision. The application of FMs in autonomous driving holds considerable promise. For example, they can contribute to enhance scene understanding and reasoning. By pre-training on rich linguistic and visual data, FMs can understand and interpret various elements in a driving scene, and provide cognitive reasoning to give linguistic and action commands for driving decisions and planning. Furthermore, FMs can augment data based on its understanding of driving scenarios to provide feasible scenes of those rare occurrences in the long tail distribution that are unlikely to be encountered during routine driving and data collection. The enhancement can subsequently lead to the improvement in the accuracy and reliability of autonomous driving systems. Another testament to the potential of FMs applications lies in the development of World Models, exemplified by the DREAMER series, which showcase the ability to comprehend physical laws and dynamics. Learning from massive data under the paradigm of self-supervised learning, World Model can generate unseen yet plausible driving environment, facilitating the enhancement in the prediction of road users behavior and the off-line training of driving strategies. In this paper, we synthesize the applications and future trends of FMs in autonomous driving. By utilizing the powerful capabilities of FMs, we strive to tackle the potential issues stemming from the long-tail distribution in autonomous driving, consequently advancing overall safety in this domain.
@@ 1 Introduction
A Foundation Model (FM) is commonly perceived as a large-scale machine learning model trained on diverse data, capable of being applied to various downstream tasks, which might not necessarily be directly related to its original training objective. The term is coined by The Stanford Institute for Human-Centered Artificial Intelligence’s (HAI) Center in August 2021 as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks” [1]. Examples of FMs include BERT[2] and the GPT (Generative Pre-Trained Transformer) series[3] in the field of Natural Language Processing (NLP), and CLIP[4] in the field of Computer Vision (CV). Most FMs are constructed based on pre-existing architectures. For example, BERT and the GPT series (language models) are based on the Transformer architecture[5], CLIP (a vision model) is founded on the Vision-Transformer[6], and DriveDreamer[7] (a world model) is built upon the Diffusion Model[8].
@@ 1.1 Emergent Abilities
Currently, the emergent abilities of FMs are mainly reflected in Large Language Models (LLMs). In Fig. 1[10] it is illustrated that as the model size, dataset size and the number of computational floats used for training increase, the loss of LLM decreases, providing support for performing large-scale model training. In Fig. 2[9], it can be seen that when the number of parameters of the model reaches a certain level, the performance of the LLMs will get a qualitative leap, showing emergent abilities in different tasks.
Emergent abilities of LLMs are well represented in In Context Learning (ICL)[11]. Context learning ability is a machine learning method to learn in a specific contextual environment. The core idea of this method is to learn from analogies[12], which enables LLMs to get excellent performance in a specific context without parameter tuning.
One particular type of ICL is Chain-of-Thought (CoT)[13]. It allows users to record the derivation process step-by-step for complex problems, such as reasoning problems, and make it available to LLMs. In this way, the LLMs can perform relatively complex reasoning tasks. Emergent abilities are commonly found in FMs, there is currently no compelling explanation for why these abilities would appear the way they do.
@@ 1.2 Pre-training
The implementation of FMs is based on transfer learning and scaling[1]. The idea of transfer learning[15, 16] is to apply the knowledge learned in one task to another. In deep learning, transfer learning is implemented in two stages, pre-training and fine-tuning. FMs are pre-trained with massive data. After obtaining the pre-trained model a specific dataset is selected for fine-tuning to adapt to different downstream tasks.
Pre-training is the foundation of FMs to obtain emergent abilities. By being pre-trained with massive data, FMs can obtain basic understanding and generative abilities. Pre-training tasks include supervised learning, self-supervised learning, etc[17]. Early pre-training relied on supervised learning, especially in computer vision. To meet the training needs of neural networks, some large-scale supervised datasets, such as ImageNet[18], were built. However, supervised learning also has some drawback, i.e., large-scale data labeling is required. With the gradual increase in the size of the model and the number of parameters, the drawbacks of supervised learning become more obvious. In natural language processing, since the degree of difficulty in labeling text is much greater than that of in labeling images, self-supervised learning is gradually favored by scholars due to its feature of not requiring labeling.
@@ 1.2.1 Self-Supervised Learning
Self-supervised learning allows learning feature representations in unlabeled data for subsequent tasks. The distinguishing feature of self-supervised learning is that they do not require manually labeled labels, but instead generate labels automatically from unlabeled data samples.
Self-supervised learning usually involves two main processes[19]: Self-supervised training phase: the model is trained to solve a designed pretext task and automatically generate pseudo labels for the pretext task based on data properties in this phase. The goal is to make the model learn the generic representation of the data.
Downstream tasks application phase: after the completion of self-supervised training, the knowledge learned by the model can be further applied to actual downstream tasks. Downstream tasks use supervised learning methods after self-supervised pre-training of the model. Downstream tasks include semantic segmentation[20], target detection[21], and sentiment analysis[22], etc. Due to the presence of self-supervised training, the generalization ability and convergence speed of the model in downstream tasks will be greatly improved.
Self-supervised learning methods can be generally categorized into three types [23]: generative-based, contrastive-based, and adversarial-based. Generative-based method: first encodes the input data using an encoder and then uses a decoder to regain the original form of the data. The model learns by minimizing the error between the two. And generative-based methods include auto-regressive models, auto-encoding models, etc[24]. Contrastive-based method: constructs positive and negative samples by pretext tasks and learns by comparing the similarity with positive and negative samples. Such methods include SimCLR[25] and others. Adversarial-based method: this method consists of two main components: a generator and a discriminator. The generator is responsible for generating fake samples, while the discriminator is used to distinguish between these fake samples and real samples[23], and a typical example is generative adversarial network[26].
@@ 1.2.2 Pretext Tasks of Self-Supervised Learning
The pretext tasks can also be referred to as self-supervised tasks as they rely on the data itself to generate labels. These tasks are designed to enable the model to learn representations that are relevant to a specific task, thereby better handling downstream tasks.
In computer vision, the design of pretext tasks according to data attributes can be summarized into four categories[19]: generation-based, context-based, free semantic label-based, and cross-modal-based. Among them, generation-based approaches mainly involve image or video generation tasks[27, 28]; context-based pretext tasks are mainly designed leveraging contextual features of images or videos, such as contextual similarity, spatial structure, temporal structure, etc.[29, 30, 31]; in the free semantic label-based pretext tasks, the network is trained leveraging automatically generated semantic labels[32]; and cross modal-based pretext tasks need to consider multiple modalities such as vision and voice[33].
In natural language processing, the most common pretext tasks include[34]: center and neighbor word prediction, next and neighbor sentence prediction, autoregressive language modeling, sentence permutation, and masked language modeling, etc. Among them, the Word2Vec[35] model uses center word prediction as a pretext task, while the BERT model uses next sentence prediction and masked language modeling as pretext tasks. These models are trained to learn the expressions of the corpus and applied to downstream tasks.
@@ 1.3 Fine-tuning
Fine-tuning refers to further training on a specific task based on an already trained model to adapt it to the specific data and requirements of the task. Typically, a model that has been pre-trained on large-scale data is used as a foundational model, and then it is fine-tuned on a specific task to improve performance. Currently, in the field of LLMs, fine-tuning methods include two main approaches: instruction tuning and alignment tuning[36].
Instruction fine-tuning is aimed to fine-tuning pre-trained models on a collection of datasets described by instructions[37]. Instruction fine-tuning is implemented by collecting or creating instances of instruction formatting and then using these instances to fine-tune the model in a supervised learning manner. Instruction fine-tuning allows LLMs to exhibit strong generalization capabilities on previously unseen tasks.
@@
In most cases, the models obtained after pre-training and fine-tuning can work well, however, some special cases may occur. In the case of LLM, for example, the trained model may appear to fabricate false information or retain biased information from the corpus. To avoid such problems, the concept of human-aligned fine-tuning has been proposed, which aims to make the model’s behavior conform to human expectations[38]. In contrast to instruction fine-tuning, this kind of alignment requires the consideration of completely different standards.
@@
The GPT family is a typical FM, and its training process also includes two stages pre-training and fine-tuning. Taking ChatGPT as an example, the pre-training process of ChatGPT uses self-supervised pre-training [3]. Given an unsupervised corpus, a standard language modeling approach is used to optimize its Maximum Likelihood Estimate(MLE). GPT uses a multi-layer transformer decoder architecture[39], resulting in a pre-trained model.
@@
The fine-tuning phase of ChatGPT can be divided into three steps[38]. In the first step, supervised fine-tuning (SFT) is performed on the obtained pre-trained model; in the second step, comparison data are collected to train the reward model (RM); and in the third step, the SFT model is fine-tuned to maximize the reward leveraging the PPO algorithm[40]. The last two steps together are Reinforcement Learning with Human Feedback (RLHF)[41].
1.4 Applications in Autonomous Driving
The ultimate goal of autonomous driving is to achieve a driving system that can completely replace human driving, and the basic criterion for evaluation is to drive like a human driver. In the past decade, autonomous driving has been rapidly developed in both academia and industry. However, in the face of uncommon or extreme scenario data in autonomous driving scenarios, the existing autonomous driving techniques have limited ability to understand and reason about such scenarios, which still greatly restricts the further development of autonoumous driving. FMs based on pre-training and fine-tuning under the paradigm unleash the great potential of AI in advanced reasoning tasks, and the use of FMs can effectively enhance the understanding and reasoning of the driving scenarios of the autonomous driving system. The research related to the realization of autonomous driving systems by empowering them with FMs has attracted great attention from academia and industry.
In this paper, we provide an overview of the application of FMs in autonomous driving in three segments. Section 2 introduces the application of language and vision FMs in enhancing autonomous driving models. Section 3 introduces the application of language and vision FMs for data augmentation. Section 4 describes the application of World Models, a class of FMs, in exploring the field of autonomous driving. Based on the overview analysis in Sections 2-4, Section 5 presents the challenges and future directions for augmenting autonomous driving with FMs.
2 Autonomous Driving Models
2.1 Human-like Driving Based on Language and Vision Models

### v2

@@ Abstract
With the development of artificial intelligence and breakthroughs in deep learning, large-scale Foundation Models (FMs), such as GPT, Sora, etc., have achieved remarkable results in many fields including natural language processing and computer vision. The application of FMs in autonomous driving holds considerable promise. For example, they can contribute to enhancing scene understanding and reasoning. By pre-training on rich linguistic and visual data, FMs can understand and interpret various elements in a driving scene, and provide cognitive reasoning to give linguistic and action instructions for driving decisions and planning. Furthermore, FMs can augment data based on the understanding of driving scenarios to provide feasible scenes of those rare occurrences in the long tail distribution that are unlikely to be encountered during routine driving and data collection. The enhancement can subsequently lead to improvement in the accuracy and reliability of autonomous driving systems. Another testament to the potential of FMs’ applications lies in World Models, exemplified by the DREAMER series, which showcases the ability to comprehend physical laws and dynamics. Learning from massive data under the paradigm of self-supervised learning, World Model can generate unseen yet plausible driving environments, facilitating the enhancement in the prediction of road users’ behaviors and the off-line training of driving strategies. In this paper, we synthesize the applications and future trends of FMs in autonomous driving. By utilizing the powerful capabilities of FMs, we strive to tackle the potential issues stemming from the long-tail distribution in autonomous driving, consequently advancing overall safety in this domain.
@@ 1 Introduction
Autonomous driving, as one of the most challenging tasks in artificial intelligence, has received considerable attention. The conventional autonomous driving system adopts a modular development strategy[yurtsever2020survey, grigorescu2020survey], whereby perception, prediction, and planning are developed separately and integrated into the vehicle. However, the information transmitted between modules is limited, and there is missing information. Furthermore, there are cumulative errors in the propagation process, and the computational efficiency of the modular transmission is relatively low. These factors collectively result in poor model performance. To further reduce the error and improve the computational efficiency, in recent years, researchers have attempted to train the model in an end-to-end manner[96, chib2023recent]. End-to-end means that the model takes inputs directly from the sensor data and then outputs control decisions for the vehicle directly. While some progress has been made, the models still mainly rely on supervised learning with manually labeled data. Due to the ever-changing driving scenarios in the real world, it is challenging to cover all potential situations with only limited labeled data. This results in a model with poor generalization ability, which makes it difficult to adapt to the complex and changeable real-world driving corner cases.
In recent years, the emergence of Foundation Models (FMs) has provided new ideas to address this gap. A Foundation Model is commonly perceived as a large-scale machine learning model trained on diverse data, capable of being applied to various downstream tasks, which might not necessarily be directly related to its original training objective. The term was coined by Stanford University in August 2021 as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks” [1]. Examples of FMs include BERT[2] and the GPT-4[43] in Natural Language Processing (NLP), and Sora[videoworldsimulators2024] in Computer Vision (CV). Most FMs are constructed based on pre-existing architectures. For example, BERT and GPT-4 are based on the Transformer[5], and Sora is founded on the Diffusion Transformer[peebles2023scalable].
Different from traditional deep learning, FMs can learn directly from massive unlabeled data (e.g., videos, images, natural language, etc.) through self-supervised pre-training, thereby acquiring stronger generalization ability and emergent abilities (thought to have already appeared in Large Language Models). Based on this, after fine-tuning with a small amount of supervised data, FMs can be rapidly adapted and migrated to downstream tasks such as autonomous driving. With the strong comprehension, inference, and generalization ability imparted by self-supervised pre-training, FMs are expected to break the bottleneck of traditional models, enabling the autonomous driving system to better understand and adapt to complex traffic environments, thus providing a safer and more reliable autonomous driving experience.
@@ 1.1 Emergent Abilities
Currently, the emergent abilities of FMs are mainly reflected in LLMs. In Fig. 1[10] it is illustrated that as the model size, dataset size, and the number of computational floats used for training increase, the loss of LLM decreases, providing support for performing large-scale model training. Fig. 2[9] shows that when the amount of parameters of the model reaches a certain level, the capabilities of LLMs will get a qualitative leap, showing emergent abilities in different tasks.
The emergent abilities of LLMs are well represented in In Context Learning (ICL)[9, 11], which, strictly speaking, can be regarded as a subclass of prompt tuning. Context learning ability is the capability of LLMs to learn in a specific contextual environment. The main idea is to learn from analogies[12]. ICL or prompt learning enables LLMs to get excellent performance in a specific context without parameter tuning.
One particular type of ICL is Chain-of-Thought (CoT). Users can break down complex problems into a series of reasoning steps as input to LLM. In this way, LLM can perform complex reasoning tasks[13]. Emergent abilities are commonly found in LLMs, there is currently no compelling explanation for why these abilities would appear the way they do.
@@ 1.2 Pre-training
The implementation of FMs is based on transfer learning and scaling[1]. The idea of transfer learning[15, 16] is to apply the knowledge learned in one mission to another. In deep learning, transfer learning is implemented in two stages, pre-training and fine-tuning. FMs are pre-trained with massive data. After obtaining the pre-trained model a specific dataset is selected for fine-tuning to adapt to different downstream tasks.
Pre-training is the foundation for FMs to obtain emergent abilities. By being pre-trained with massive data, FMs can obtain basic understanding and generative capability. Pre-training tasks include supervised learning (SL), self-supervised learning (SSL), etc[17]. Early pre-training relied on supervised learning, especially in computer vision. To meet the training needs of neural networks, some large-scale supervised datasets, such as ImageNet[18], were built. However, supervised learning also has some drawbacks, i.e., large-scale data labeling is required. With the gradual increase in the size of the model and the amount of parameters, the drawbacks of supervised learning become more obvious. In NLP, since the degree of difficulty in labeling text is much greater than that of labeling images, SSL is gradually favored by scholars due to its feature of not requiring labeling.
@@ 1.2.1 Self-Supervised Learning
Self-supervised learning allows learning feature representations in unlabeled data for subsequent tasks. The distinguishing feature of SSL is that they do not require manually labeled labels, but instead generate labels automatically from unlabeled data samples.
SSL usually involves two main processes[19]: Self-supervised training phase: the model is trained to solve a designed pretext task and automatically generates pseudo labels based on data properties in this phase. It is designed to allow the model to learn the generic representation of the data. Downstream tasks application phase: after the self-supervised training, the knowledge learned by the model can be further used for actual downstream tasks. Downstream tasks use supervised learning methods, which include semantic segmentation[20], target detection[21], and sentiment analysis[22], etc. Due to self-supervised training, the generalization ability and convergence speed of the model in downstream tasks will be greatly improved.
SSL methods generally fall into three categories [23]: generative-based, contrastive-based, and adversarial-based. Generative-based method: first encodes the input data using an encoder and then uses a decoder to regain the original form of the data. The model learns by minimizing the error. And generative-based methods include auto-regressive models, auto-encoding models, etc[24]. Contrastive-based method: constructs positive and negative samples by pretext tasks and learns by comparing the similarity with positive and negative samples. Such methods include SimCLR[25] and others. Adversarial-based method: this method consists of a generator and a discriminator. The generator is responsible for generating fake samples, while the discriminator is adapted to distinguishing between these fake samples and real samples[23], and a typical example is generative adversarial network[26].
@@ 1.2.2 Pretext Tasks of Self-Supervised Learning
The pretext tasks can also be referred to as self-supervised tasks as they rely on the data itself to generate labels. These tasks are designed to make the model learn representations that are relevant to a specific task, thereby better handling downstream tasks.
In computer vision, the method of designing pretext tasks according to data attributes includes four main categories[19]: generation-based, context-based, free semantic label-based, and cross-modal-based. Among them, generation-based approaches mainly involve image or video generation tasks[27, 28]; context-based pretext tasks are mainly designed leveraging contextual features of images or videos, such as contextual similarity, spatial structure, temporal structure, etc.[29, 30, 31]; in the free semantic label-based pretext tasks, the network is trained leveraging automatically generated semantic labels[32]; and cross modal-based pretext tasks need to consider multiple modalities such as vision and voice[33].
In natural language processing, the most common pretext tasks include[34]: center and neighbor word prediction, next and neighbor sentence prediction, autoregressive language modeling, sentence permutation, masked language modeling, etc. Among them, the Word2Vec[35] model uses center word prediction as a pretext task, while the BERT model uses next sentence prediction and masked language modeling as pretext tasks. These models are trained to learn the expressions of the corpus and applied to downstream tasks.
@@ 1.3 Fine-tuning
Fine-tuning is the process of further training on a specific task based on an already trained model, to adapt it to the specific data and requirements of the task. Typically, a model that has been pre-trained on large-scale data is used as a foundational model, and then it is fine-tuned on a specific task to improve performance. Currently, in the field of LLMs, fine-tuning methods include two main approaches: instruction tuning and alignment tuning[36].
@@
Instruction fine-tuning aims at fine-tuning pre-trained models on a collection of datasets described by instructions[37]. Instruction fine-tuning generally includes two phases. First, instances of instruction formatting need to be collected or created. And then these instances are used to fine-tune the model. Instruction fine-tuning allows LLMs to exhibit strong generalization ability on previously unseen tasks. The models obtained after pre-training and fine-tuning can work well in most cases, however, some special cases may occur. In the case of LLM, for example, the trained model may appear to fabricate false information or retain biased information from the corpus. To avoid such problems, the concept of human-aligned fine-tuning was proposed. The goal is to make the model’s behaviors conform to human expectations[38]. In contrast to instruction fine-tuning, this kind of alignment requires the consideration of completely different standards.
@@
The GPT family is a typical FM, and its training process also includes pre-training and fine-tuning. Taking ChatGPT as an example, the pre-training process of ChatGPT uses self-supervised pre-training [3]. Given an unsupervised corpus, a standard language modeling approach is used to optimize its Maximum Likelihood Estimate (MLE). GPT uses a multi-layer transformer decoder architecture[39], resulting in a pre-trained model.
@@
The fine-tuning phase of ChatGPT consists of the following three steps[38]. Firstly, supervised fine-tuning (SFT) is performed on the obtained pre-trained model; secondly, comparison data are collected to train the reward model; and thirdly, the SFT model is fine-tuned to maximize the reward leveraging the PPO algorithm[40]. The last two steps together are Reinforcement Learning with Human Feedback (RLHF)[41].
1.4 Abilities of Foundation Models in Autonomous Driving
The ultimate goal of autonomous driving is to realize a driving system that can completely replace human driving, and the basic criterion for evaluation is to drive like a human driver, which puts forward very high requirements on the reasoning ability of autonomous driving models. We can see that FMs based on large-scale data learning have powerful reasoning and generalization ability, which have great potential in autonomous driving. FMs can be used to enhance scenarios understanding, give language-guided commands, and generate driving actions in empowering autonomous driving. In addition, FMs can be augmented with powerful generative capability for data augmentation, including extending existing autonomous driving datasets and directly generating driving scenarios. In particular, World Models (a type of FM) can learn the inner workings of the physical world and predict future driving scenarios, which is of significant importance for autonomous driving.
Consequently, it was deemed appropriate to conduct a comprehensive review of the applications of FMs in autonomous driving. This paper provides that review.
  • In Section 2, a brief overview of the latest supervised end-to-end autonomous driving is provided, to offer the reader a better background understanding.
  • Section 3 reviews the applications of language and vision FMs in enhancing autonomous driving.
  • Section 4 reviews the applications of World Models in the exploration of the field of autonomous driving.
  • Section 5 reviews the applications of FMs in data augmentation.
Building on the preceding overview, Section 6 presents the challenges and future directions for enhancing autonomous driving with FMs.
2 Supervised End-to-end Autonomous Driving
The research idea of “pre-training + fine-tuning” in autonomous driving research has not only appeared after the introduction of large models but has been researched for a long time. To use a more familiar term, it is end-to-end autonomous driving. In the past few years, some scholars have already optimized Pretraining Backbone in various ways, including the transformer architecture and self-supervised learning methods——Note that we define Pretraining Backbone here refers to a model that transforms each modal input into a usable feature representation for downstream tasks (such as target detection, trajectory prediction, decision planning, etc.). Many research attempts have also been made to develop end-to-end frameworks based on the transformer architecture, with excellent results. Hence, to summarize the application of the underlying models in autonomous driving more comprehensively, we believe that it is necessary to introduce the research related to end-to-end autonomous driving based on Pretraining Backbone. In this section, we summarize the latest research on Pretraining Backbone with end-to-end autonomous driving solutions. The pipeline for such methods is briefly illustrated in Fig. 3.
Figure 3: The pipeline diagram for the supervised end-to-end autonomous driving system with a Pretraining Backbone. Multi-modal sensing information is input to the Pretraining Backbone to extract features, after which it enters into the framework of autonomous driving algorithms built by various methods to realize tasks, such as planning/control, to accomplish end-to-end autonomous driving tasks.
2.1 Pretraining Backbone
In end-to-end modeling, feature extraction of low-level information from raw data determines the potential of subsequent model performance to a certain extent, and an excellent Pretraining Backbone can endow the model with more powerful feature learning capability.
Pretraining convolutional networks such as ResNet[79], and VGG[80] are the most widely used backbone for visual feature extraction in end-to-end models. These pretraining networks are often trained to leverage target detection or segmentation as the task to extract generalized feature information, and their competitive performances have been verified in many works. ViT[6] first applied the transformer architecture to image processing and achieved excellent classification results. Transformer has the advantage of optimized algorithms for handling large-scale data with its simpler architecture and faster inference speed. The self-attention mechanism is very suitable for processing time series data, there is the possibility of modeling and predicting the temporal motion trajectories of objects in the environment, and is conducive to the fusion of heterogeneous data from multiple sources, such as LiDAR point clouds, images, maps, etc.
Another class of Pretraining Backbone, represented by LSS[81], BEVDet[82], BEVformer[83], BEVerse[84], etc., expanded the usability by extracting the images captured by the surround-view camera and converting them to Bird’s-Eye View (BEV) features through model learning, indexing the local image features from the 2D-viewpoints to the 3D-space. In recent years, BEV has attracted extensive interest due to its ability to describe the driving scene more accurately, and the research of leveraging BEV features such as Pretraining Backbone output is not limited to camera, and the extraction and fusion of multi-modal sensing BEV features represented by BEVfusion[85] has further provided a wider vision for autonomous driving systems. However, it should be pointed out that although the transformer architecture brings great performance enhancement, these Backbone still constructs pre-trained models with supervised learning methods, which rely on massive labeled data, and the data quality also greatly affects the final result of the model.
In both camera and point cloud processing domains, some works implement the Pretraining Backbone by unsupervised or self-supervised learning methods. Wu et al.[86] proposed the PPGeo model, which uses a large number of unlabeled driving videos to accomplish the pre-training of the visual coder in two stages, and can be adapted to different downstream end-to-end autonomous driving tasks. Sautier et al.[87] proposed BEVContrast for self-supervision of 3D Backbone on automotive LiDAR point clouds, which defines contrasts at the level of 2D Cells in the BEV plane, retaining the simplicity as in PointContrast[88] while maintaining good performance in downstream driving tasks. Especially, while the self-supervised learning approach of ”masking + reduction” is also considered to be an effective way of modeling the world, Yang et al.[yang2023unipad] proposed UniPAD, which is implemented based on self-supervised learning methods for MAE and 3D rendering. A portion of these multi-modal data is randomly keyed out to be masked and transformed into voxel space, where RGB or depth prediction results are generated by rendering techniques in such a 3D space, and the rest of the original images are used as the generated data for supervised learning. The flexibility of the approach enables good integration into both 2D and 3D frameworks, and downstream tasks such as depth estimation, target detection, segmentation, and many others fine-tuned and trained on the model perform superiorly.
2.2 Supervised End-to-end Autonomous Driving Models
Early work on end-to-end autonomous driving modeling was mainly based on various types of deep neural networks, which were constructed through imitation learning[89, 90, 91, 92, hu2022st] or reinforcement learning[93, 94, 95] methods. The work of Chen et al.[96] analyzed the key challenges facing end-to-end autonomous driving from a methodological perspective, pointing out the future trend of empowering end-to-end autonomous driving with fundamental models such as transformer. Some scholars have tried to build an end-to-end autonomous driving system with transformer and got competitive results. For instance, there have been Transfuser[97, 98], NEAT[99], Scene Transformer[100], PlanT[101], Gatformer[102], FusionAD[103], UniAD[104], VAD[jiang2023vad], GenAD[zheng2024genad] and a host of end-to-end frameworks developed based on transformer architecture.
Chitta et al.[97, 98] proposed Transfuser, which takes RGB images and BEV views from LiDAR as inputs, uses multiple transformers to fuse the feature maps, and predicts the trajectory points for the next four steps through a single-layer GRU network, followed by longitudinal and transverse PIDs to control the vehicle operation. NEAT[99] further mapped the BEV scene to trajectory points and semantics information, then used an intermediate attention map to compress high-dimensional image features, which allows the model to focus on driving-relevant regions and ignore driving task-irrelevant information. PlanT proposed by Renz et al.[101] used simple object-level representations (vehicles, roads) as inputs to the transformer encoder, and used speed prediction of surrounding vehicles as a secondary task to predict future waypoint trajectories. UniAD proposed by Hu et al.[104] enhanced the design of the decoder and achieved the integration of the full stack of autonomous driving tasks into a unified framework to improve autonomous driving performance, although still relying on different sub-networks for each task. This work also won the CVPR 2023 Best Paper Award, which shows the academic recognition of the end-to-end autonomous driving paradigm. However, these models often require intensive computation. For this reason, Jiang et al.[jiang2023vad] proposed a method to fully vectorize the driving scenarios and learn the instance-level structural information to improve computing efficiency. In contrast to the previous modular end-to-end planning, Zheng et al.[zheng2024genad] propose a generative end-to-end, modeling autonomous driving as a trajectory generation.
Moreover, Drive Anywhere proposed by Wang et al.[105] not only realizes end-to-end multi-modal autonomous driving but also combines with LLM to be able to provide driving decisions based on representations that can be queried through images and texts. Dong et al.[106] generated image-based action commands and explanations by building a feature extraction model based on transformer. Jin et al.[107] proposed the ADAPT model to directly output vehicle control signals with inference language descriptions through an end-to-end model. This is the first driving action captioning architecture based on an action-aware transformer. It accomplished the driving control task while adding natural language narratives to guide the decision-making and action process of the autonomous driving control module. It also helped the user to get the vehicle’s state and the surrounding environment at all times and to better understand the basis of the actions taken by the autonomous driving system, which improved the interpretability of the decision-making. This provides a glimpse of the potential of the transformer architecture to improve the interpretability of end-to-end driving decisions.
3 Human-like Driving Based on Language and Vision Models